How do I identify the solution of 4x+2y=0 and y=5-3x? I was able to get as far as -10, but from there I am completely lost.
Answer by Math_Boss(45)

Substitute y=(5-3x)
4x+2(5-3x)=0
4x+10-6x=0
-2x=-10
x=5
y=5-3x5=-10
Ans:(5,-10)